5 over 7 divided 9 over14=w+1....solve w
Answer by checkley75(3666)

(5/7)/(9/14)=w+1 dividing a fraction by a fraction requires the inversion of the lower fraction and then multiplying the two fractions. thus:
5/7*14/9=w+1
70/63=w+1
10/9=w+1 now cross multiply and we get:
9(w+1)=10
9w+9=10
9w=10-9
9w=1
w=1/9